OpenStudy (anonymous):

1/0.11 +12(6.12-0.11)^2 use the correct order of operations to make the following calculations

OpenStudy (arindameducationusc):

you there?????

OpenStudy (anonymous):

yes

OpenStudy (arindameducationusc):

6.12-.11=6.01 now (6.01)^2=36.12

OpenStudy (arindameducationusc):

36.12*12=433.44 and 1/0.11=9.0909

OpenStudy (arindameducationusc):

now add433.44+9.0909=442.5209 DONE!

OpenStudy (arindameducationusc):

If it was helpful dont forget to medal and fan. Thank you

OpenStudy (anonymous):

Thank you

OpenStudy (arindameducationusc):

Np